An aqueous solution of sodium fluoride is slowly added to a water sample that contains barium ion (2.75×10-2M ) and calcium ion (7.30×10-2M ). What is the remaining concentration of the first ion to precipitate when the second ion begins to precipitate?
Answer #1

We must know which ion will precipitate first. To find out this, we need to calculate minimum concentration of Fluoride ions required to start precipitation of BaF 2 and CaF2 .

Consider dissociation of BaF2 in water , BaF 2 (s) phpv6rcS4.png Ba 2+ (aq) + 2 F - (aq)

For above reaction, Ksp = [ Ba 2+ ][F - ] 2 = 1.0 \times 10 -06

\therefore [F - ] minimum = ( 1.0 \times 10 -06 / [ Ba 2+ ] ) 1/2

[F - ] minimum = ( 1.0 \times 10 -06 / 2.75 \times 10 -02 ) 1/2

[F - ] minimum = 6.030 \times 10 -03 M

Consider dissociation of CaF2 in water , CaF 2 (s) phpv6rcS4.pngCa 2+ (aq) + 2 F - (aq)

For above reaction, Ksp = [ Ca 2+ ][F - ] 2 = 5.3 \times 10 -09

\therefore [F - ] minimum = ( 5.3 \times 10 -09 / [ Ca 2+ ] ) 1/2

[F - ] minimum = ( 5.3 \times 10 -09 / 7.30 \times 10 -02 ) 1/2

[F - ] minimum =2.694 \times 10 -04 M

From above values , it is clear that Calcium fluoride will precipitate out first.

We have to find out [ Ca 2+ ] when BaF2 starts precipitates out.

When BaF2 starts precipitates out , [F - ] is 6.030 \times 10 -03 M ( \because [F - ] minimum = 6.030 \times 10 -03 M)

\therefore [ Ca 2+ ] = K sp / [F - ] = 5.3 \times 10 -09 /  6.030 \times 10 -03 = 8.79 \times 10 -07 M .

ANSWER : [ Ca 2+ ] when BaF2 starts precipitates out is 8.79 \times 10 -07 M .